OpenStudy (anonymous):

A cylindrical water glass has a radius of 7 centimeters and a height of 14 centimeters. What is its volume rounded to the nearest hundredth?

OpenStudy (anonymous):

4,310.27 cm3 1,436.76 cm3 2,155.13 cm3 718.38 cm3

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (goldphenoix):

Do you know the formula for finding the volume of a cylinder?

OpenStudy (anonymous):

v=1/3bh

OpenStudy (anonymous):

idk if that right

OpenStudy (goldphenoix):

Nope. The formula for finding the volume of a cylinder is: \[\large \large \pi \times r^2 \times h\]

OpenStudy (goldphenoix):

Now plug it in. What do you get?

OpenStudy (anonymous):

2154.04

OpenStudy (goldphenoix):

Hmm. Did you use 3.14, or the pi button?

OpenStudy (goldphenoix):

It would be better if you use the pi button to get an accurate number.

OpenStudy (anonymous):

yes

OpenStudy (goldphenoix):

Then again, you probably know your answer.

OpenStudy (goldphenoix):

So the answer is?

OpenStudy (anonymous):

2154.04

OpenStudy (anonymous):

idk

OpenStudy (goldphenoix):

I mean, which answer from the multiple choice is reasonable?

OpenStudy (anonymous):

idk

OpenStudy (goldphenoix):

Hmm... Which number from the multiple choices is the closest to 2154.04

OpenStudy (anonymous):

c

OpenStudy (goldphenoix):

Good job! So the answer is C. I hope this help. :)

OpenStudy (anonymous):

thanks